# The Cooperative Network Architecture: Learning Structured Networks as Representation of Sensory Patterns

### v3

@@ 5.3 Reconstructing Figures
Figure 7: Part A: Input images of diagonal lines with different numbers of missing pixels, corresponding feature activations (middle column), and final output neuron activations. Activation bias $b^{(S2)}=0.5$, attenuation coefficient $\gamma=0.6+0.2t$. Part B: Diagonal line with $6$ missing pixels in the center (shown on the left), the corresponding output activity (in the middle), and the membrane potential $\boldsymbol{a}^{(S2)}$ displaying the “activation probabilities” (on the right) after the first time step.
@@
If a fragmet $\mathcal{F}_{t}(n_{i})$ contains a substantial number of (active) neurons, an inactive neuron $n_{i}$ can receive significant support, leading it to switch on and encouraging figure reconstruction ($\mathcal{F}_{t}(n_{i})\rightarrow\mathcal{F}^{*}_{t}(n_{i})$). Figure 7A shows how inactive neurons activate and thus demonstrates that nets can deal with occluded patterns. This reconstruction is reliable for up to $3$ missing pixels (see Appendix G). In some cases, as in this Figure, the line is fully reconstructed for up to $6$ removed pixels and partly reconstructed for $7$ pixels.
Membrane Potential Represents Expected Feature Activation.
@@
In the presence of missing pattern elements, the membrane potential map $\boldsymbol{a}^{(S2)}$ is higher at spatial locations where line features are observed and lower where they are not observed but expected (as shown in Figure 7B). Depending on the context, it may be desirable either to reconstruct the missing line segment – effectively bridging the gap between discontinuous features – or to preserve the separation between distinct line elements. This behavior can be modulated by adjusting the hyperparameter $b^{(S2)}$ in eq. 18, which can govern the balance between feature completion and preservation of discontinuity, thereby allowing the model to selectively favor line reconstruction or separation.
5.4 Comparison to an Autoencoder
@@
Figure 8: Comparison of the outputs generated by the proposed CNA model and an autoencoder. Each row presents two samples, $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, alongside their respective feature activations and resulting representations (once for the CNA model and once for the autoencoder). Rows $1$ to $3$ illustrate how both models encode patterns that were not present in the training dataset. Rows $4$ and $5$ compare the models’ performance in denoising tasks, while row $6$ evaluates their ability to reconstruct lines.
@@
Figure 8 compares performance between the proposed CNA model and an autoencoder (for model details, refer to Appendix F). While the autoencoder successfully learns to accurately represent line structures observed in the training data – achieving a mean square error (MSE) of less than $1\times 10^{-3}$ between the input and the reconstruction by the decoder – it struggles with generalization to data not observed during training. For instance, rows $1$ to $3$ demonstrate that the CNA model is capable of representing previously unseen structures, while, in contrast, the autoencoder fails to capture such out-of-distribution patterns.
Rows $4$ and $5$ highlight the autoencoder’s ability to represent line structures with up to 11.5% noise introduced. However, as the noise level increases, its performance decreases significantly. The CNA model, by contrast, remains robust under much higher noise levels, tolerating up to $59\%$ noise while still producing high-quality representations. Metrics on noise reduction, precision, and recall, presented in Appendix G, further substantiate the CNA model’s superior performance.
@@
In row $6$, the models’ capabilities for reconstructing missing line segments are compared. Both models can reconstruct the removed pixels, but the autoencoder introduces additional noise in regions other than the line center, where the pixels were removed. This observation is confirmed by the metrics in Appendix G: The autoencoder is slightly superior in reconstructing the removed pixels, but this comes at the cost of increased noise elsewhere, leading to lower overall precision and recall for figure reconstruction.
@@ 6 Relation to Alternative Concepts
To grasp the significance of the approach, it is important to compare it to alternative approaches.
@@ 6.1 Associative Memory
Associative memory (AM) (Amari, 1972; Hopfield, 1982) and the model presented in this work share the objective of retrieving previously encountered patterns, both utilize Hebbian synaptic plasticity for encoding information. However, fundamental differences exist in their structural and functional organization. In classical AM, neurons exhibit an all-to-all connectivity pattern, leading to the representation of structures as monolithic wholes, whereas our model constrains connectivity within learned net fragments. This restriction enables a more modular representation of complex patterns by dynamically assembling net fragments into structured nets, thereby allowing the formation of representations that extend beyond the training data distribution.
These differences arise from distinct learning mechanisms. In AM, a neural pattern is stored in a single event through the uniform strengthening of connections between co-activated neurons. Our approach, however, capitalizes on the observation that a large number of non-repeating patterns (as found in natural images) contain small recurring sub-patterns, such as texture patches. These recurring elements are captured within net fragments. The learning rule defined in eq. 20 selectively reinforces connections corresponding to the most relevant sub-patterns. Specifically, frequent sub-patterns exhibit a consistently positive average correlation $\rho$, which leads to the progressive increase of their associated synaptic weights. In contrast, rare or non-informative patterns exhibit low or even negative average correlations, causing their associated weights to decay toward zero. This Hebbian mechanism effectively filters out noise and ensures that only the most salient and statistically dominant sub-patterns are preserved in the learned representation.
A fundamental limitation of AM is the interference between stored patterns unless they are statistically independent (i.e., orthogonal), meaning that minimal neuron-wise overlap exists. Sparse coding, as proposed by Tsodyks and Feigel’man (2007) or Palm (2013), mitigates interference by activating only a small subset of neurons. Our model offers a more intrinsic solution by representing each input feature using multiple alternative neurons. These neurons possess identical receptive fields (as observed by Cossell et al. (2015)), but can exhibit different excitatory outputs. This redundancy enables overlapping net fragments to reduce neuron-wise interference by utilizing distinct alternative neurons.
@@ 6.2 Boltzmann Machine
Key aspects of our model become clearer when compared to the Boltzmann Machine (BM) introduced by Hinton and Sejnowski (1983). Both models allow recurrent connections (as also proposed in Mumford (2002)), but they also exhibit significant differences. Nevertheless, as our focus is limited to one area with net fragments ($S2$) and is not the full object detection architecture as proposed by von der Malsburg et al. (2022), we highlight only select differences here.
The BM makes a clear distinction between visible and hidden (“latent”) neurons. In the CNA this distinction is softened: it posits neurons that are activated by sensory input, but of which the majority are quickly silenced by attenuation in favor of those neurons that support each other through previously learned connections. A winning alternative neuron can be interpreted as the hypothesis not only that the neuron has detected its trigger pattern in its afferents but also that the neuron has connections to a number of other active neurons and is consistent with them in light of previous experience.
In both BM and our model, the activity of hidden, or in our case alternative, neurons, is controlled by the same type of dynamic equations, with the difference being that we do not insist on symmetric connections and therefore do not derive an energy function (see, however, the remark made in Hinton and Sejnowski (1983), that given sufficiently redundant support of active neurons, the absence of up to half of the connections due to lack of symmetry does not jeopardize the stability of a state). However, even without an energy function, convergence is not an issue, as alternative neurons undergo winner-take-all competition.
@@ 6.3 Artificial Neural Networks (ANN)
Artificial neural networks (ANNs) rely on various architectures and learning techniques, including multilayer perceptrons (MLPs) (Prince, 2023), transformers (Vaswani et al., 2017), and autoencoders (Hinton and Salakhutdinov, 2006). These systems process sensory input through multiple layers to generate a neuronal representation. A defining characteristic of ANN systems is their strict feedforward architecture, where connections from input to representation and from representation to output are trained via backpropagation of error (Rosenblatt, 1962, p. 292).
In contrast, the neuronal representation within the CNA consists of nets that activate as a whole, with individual neurons depending on the activation of others in the same net. Thus, while ANNs rely on a global structuring of neuronal activations, the CNA relies on a local self-organizing process.
Moreover, ANNs rely on hierarchical layers where neurons respond to features of increasing complexity. The CNA replaces this hierarchy with a single layer, where net fragments of different sizes capture features of varying complexity. These fragments interconnect to form larger nets that represent entire objects, naturally solving the binding problem, whereas ANNs do not encode explicit relationships between features, leading to ambiguities such as hallucinated feature conjunctions.
Furthermore, input representation in the CNA emerges from statistical regularities in the input, making it independent of specific tasks. In contrast, ANNs bias feature extraction to serve predefined training objectives, which limits their adaptability when retraining for different functions.
@@ 7 Discussion and Outlook
The CNA presented here establishes a novel approach for constructing coherent nets as representations of input data. This framework enables the encoding of complex patterns through the composability of previously learned sub-patterns (net fragments). Compared to existing neural coding frameworks, CNA offers distinct advantages, particularly in terms of robustness – demonstrated through its ability to filter noise and reconstruct figures – and its capacity to generalize to out-of-domain data, as evidenced by its capability to represent previously unseen objects.
For future research, we outline two key directions: First, an essential next step is the extension of CNA to natural image domains. This would allow net fragments to capture and encode texture information encountered in real-world scenarios, thus realizing the ambition expressed in Figure 1. Second, the current implementation of CNA is restricted to a single processing area with net fragments. Following the theoretical framework proposed by von der Malsburg et al. (2022), the architecture could be expanded to incorporate multiple areas, thereby enabling object recognition while achieving invariance to position, scale, and in-plane rotation. Object recognition could be realized using structure-sensitive, homeomorphic mappings between nets (von der Malsburg, 1981; Anderson and van Essen, 1987; Olshausen et al., 1995; Wiskott et al., 1997; Wolfrum et al., 2008; Memisevic and Hinton, 2010; Arathorn et al., 2013), which would facilitate the mapping of input representations onto a more stable schematic representation of an object class or an exemplar of the object to be recognized. The adoption of homeomorphic mappings, as opposed to alternative methods, preserves the core principle of nets and net fragments within and across multiple processing areas, thereby maintaining structural coherence in representation.
The proposed neural coding framework not only advances representational robustness but also aligns with principles of self-organization observed in cortical processing. Its ability to learn and flexibly recombine net fragments without explicit supervision suggests a pathway toward more adaptive and interpretable neural representations. Future extensions to multi-area architectures could bridge the gap between bottom-up feature learning and top-down predictive coding. These developments position CNA as a foundational step toward more structured and generalizable neural computation.

### v4

@@ 4.5 Reconstructing Figures
Figure 7: Recovery from occlusion. Part A: Input images of diagonal lines with different numbers of missing pixels, corresponding feature activations (middle column), and final output neuron activations. Activation bias $b^{(S2)}=0.5$, attenuation coefficient $\gamma=0.6+0.2t$. Part B: A diagonal line with 666 missing pixels in the center (shown on the left), the corresponding output activity (in the middle), and the pre-synaptic activity $\boldsymbol{a}^{(S2)}$ displaying the “activation probabilities” (on the right) after the first time step.
@@
If a net fragment $\mathcal{F}_{t}(n_{i})$ contains a substantial number of (active) neurons, an inactive neuron $n_{i}$ can receive significant support, leading it to switch on and encouraging figure reconstruction ($\mathcal{F}_{t}(n_{i})\rightarrow\mathcal{F}^{*}_{t}(n_{i})$). Figure 7A shows how inactive neurons activate and thus demonstrates that nets can deal with occluded patterns. This reconstruction is reliable for up to 333 missing pixels (see Appendix G). In some cases, as in this figure, the line is fully reconstructed for up to 666 removed pixels and partly reconstructed for 777 pixels.
Pre-Synaptic Activity Represents Expected Feature Activation.
@@
In the presence of missing pattern elements, the pre-synaptic activity map $\boldsymbol{a}^{(S2)}$ is higher at spatial locations where line features are observed and lower where they are not observed but expected (as shown in Figure 7B). Depending on the context, it may be desirable either to reconstruct the missing line segment – effectively bridging the gap between discontinuous features – or to preserve the separation between distinct line elements. This behavior can be modulated by adjusting the hyperparameter $b^{(S2)}$ in Eq. 29, which can govern the balance between feature completion and preservation of discontinuity, thereby allowing the model to selectively favor line reconstruction or separation. In future work, incorporating additional areas that maintain object-associated representations and act as top-down hypotheses could enable this selectivity to be learned rather than manually tuned.
4.6 Comparison to Autoencoders
@@
Figure 8: Comparison of the outputs generated by the proposed CNA model and an autoencoder. Each row presents two samples, $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, alongside their respective feature activations and resulting representations (once for the CNA model and once for the autoencoder). Rows 111333 illustrate how both models encode patterns that were not present in the training dataset. Rows 444555 let one compare the models’ performance in denoising tasks, while row 666 displays their ability to reconstruct lines.
@@
Figure 8 compares performance between the proposed CNA model and an autoencoder (for model and training details, refer to Appendix F). While the autoencoder successfully learns to accurately represent line structures observed in the training data – achieving a mean square error (MSE) of less than $1\times 10^{-3}$ between the input and the reconstruction by the decoder – it struggles with generalization to data not observed during training. For instance, rows 111333 demonstrate that the CNA model is capable of representing previously unseen structures, while, in contrast, the autoencoder fails to capture such out-of-distribution patterns.
[…]
@@
In row 666, the models’ capabilities for reconstructing missing line segments are compared. Both models can reconstruct the removed pixels, but the autoencoder introduces additional noise in regions other than the line center, where the pixels were removed. This observation is confirmed by the metrics in Appendix G: The autoencoder is slightly superior in reconstructing the removed pixels, but this comes at the cost of increased noise elsewhere, leading to lower overall precision and recall for figure reconstruction.
@@ 5 Relation to Alternative Concepts
The central issue behind our model is how the structure of the environment is represented. In biology, this structure would be accessed by the senses in the form of sensory patterns — large active subsets of sensory elements (e.g., retinal ganglion cells). While global sensory patterns vary across different conditions and viewpoints, local sensory structure tends to exhibit consistency. For instance, the sensory pattern associated with a static object may appear different under varying eye positions or perspectives, yet local fragments of the pattern remain stable and coherent (we demonstrate this in our experiments by having constant straight-line patterns that are stabilized, while changing Gaussian noise is ignored). The task of learning – both in biology and in a model like ours – is to pick up durable structures from sensory patterns in spite of their fleeting nature. We proceed here with a simple assumption regarding this durable structure: There is an important class of coherent sensory patterns that are composed of a finite set of overlapping pattern fragments. We assume that these pattern fragments can be represented by the recurrent connectivity of net fragments and that this connectivity can be learned from the statistics of sensory patterns with the help of Hebbian plasticity. This is the background against which we discuss some models of neural computation currently dominating the literature.
@@ 5.1 Associative Memory
The central concept in the classic works of Hebb (1949) and Hayek (1952) is the assembly, a set of neurons supporting each other’s activity by mutual excitatory connections, connections that have been created by synaptic, “Hebbian”, plasticity. This concept has found a concrete mathematical formulation as “associative memory” (AM) (Amari, 1972; Hopfield, 1982), where stored patterns are encoded by enhancing excitatory connections between all pairs of co-active neurons.
While AM mechanisms are effective for storing and retrieving known patterns, they often fail when applied to perceptual tasks such as object recognition, where previously unseen structures must be classified based on prior experience. This limitation arises because AM tends to store complete patterns as static “snapshots.” The reason for this rigidity of AM states is their mode of recording: The excitatory connections between all pairs of neurons active at a particular moment are strengthened simultaneously, without regard to whether a connection is significant in the sense of applying to many patterns (that is, being part of a repeating pattern fragment) or insignificant, applying only to this particular pattern.
In our model, in contrast, the synaptic weights are gradually modified in response to the statistical properties of a large ensemble of sensory patterns. According to our Eq. 30, a particular synapse can only grow if the pair of neurons it connects is simultaneously active again and again (indicated by positive correlation $\boldsymbol{\rho}_{avg}$), or it decays due to negative correlation. This processing allows the system to gradually extract and store re-occurring (statistically relevant) sub-patterns rather than memorizing isolated global patterns.
A functional restriction on AM is that different stored states have to be statistically independent, that is, as sets of active neurons, they have to have low pairwise overlap. This is an awkward constraint, as sensory patterns are composed of repeating pattern fragments and thus are bound to have a large overlap. In classical AM, this constraint has been addressed just by requiring patterns to be small subsets of all neurons (Tsodyks and Feigel’man, 2007; Palm, 2013). Our model can address the small-overlap constraint in a systematic way with the help of multiple neurons (“competing neurons”) that respond to the same trigger feature but compete with each other for activity so that pairs of pattern fragments that have high feature-wise overlap can select different competing neurons, resulting in the corresponding net fragments having low neuron-wise overlap. This is made possible by the statistical mode of strengthening synapses and is not possible in AM with its one-shot learning mode.
Our model also addresses an open issue of AM: The selection of coherent sensory patterns for storage. As coherence is defined as being composed of a fixed set of overlapping pattern fragments (stored in net fragments), it is only natural to require that the learning process converges to a connectivity state in which only sets of neurons that segment the input into coherent sensory patterns can be activated stably (von der Malsburg, 2024).
The canonical formulation (Amari, 1972; Hopfield, 1982) of the AM’s state dynamic defines an “energy” that acts as a Lyapunov function. Beginning from any initial state, neurons are visited sequentially, and their state is flipped to gradually reduce the energy (simultaneous switching of the neurons would lead to non-sensical oscillations). Our model’s dynamics is not based on a Lyapunov function and a time-consuming gradual process, relying instead on an externally imposed periodic attenuation control parameter that lets the set of neurons originally activated by the sensory input collapse to a subset formed by neurons that support each other by excitatory connections.
@@ 5.2 Boltzmann Machine
Regarding Boltzmann Machines (BM), we refer to Hinton and Sejnowski (1983); Hinton et al. (1984), taking the latter as the authoritative reference. Both the BM and our model describe networks of recurrent connections between neurons, where some neurons get activated by sensory input (“visible units”), and by employing synaptic plasticity, both approaches aim to turn the system into a constraint satisfaction network that uses inner connections to stabilize the patterns suggested by the sensory input.
The BM as described in Hinton et al. (1984) and our model differ, however, in a number of ways. Inspired by statistical mechanics, the BM speaks of equilibrium macro-states, each of which is realized as an ensemble of micro-states, the probability of a micro-state being proportional to a negative exponential of a state “energy” (Boltzmann distribution), which is defined as in Amari-Hopfield’s Associative Memory as a function of neural activities and connection weights. The BM further assumes that the system compares two modes, clamped (in which the sensory neurons are determined by the input) and free running (no sensory input), and lets synaptic plasticity equalize the signal correlations of pairs of neurons observed in one of the two modes with those observed in the other. In our model, neural activity is not modeled in terms of probabilistic ensembles of states but as discrete-time trajectories over a fixed number of time steps $T$, during which the same sensory input is constantly refined. These trajectories evolve deterministically according to update rules that depend on both the input and recurrent connectivity. Importantly, the dynamics are externally modulated: At each step, a control mechanism attenuates the activity of neurons that lack sufficient recurrent support from other active neurons; see Eqs. 28 and 29. In contrast, Boltzmann Machines rely on stochastic sampling over ensembles of network states, governed by a distribution derived from an energy function. While the statistical mechanics framework underlying Boltzmann Machines is elegant and conceptually rich, it is computationally inefficient in settings where a large number of intermediate states must be generated and compared to approximate equilibrium. Our model avoids this bottleneck by converging rapidly through deterministic dynamics over a small fixed number of steps.
A necessary precondition for the feasibility of the goals of the BM and our model is the existence of a system of constraints on the sensory patterns that can be learned from a finite number of such patterns and which then suffice to stabilize all future patterns. Given the bias-variance dilemma (Geman et al., 1992) and the no-free-lunch theorems (Wolpert, 1996), a learning system’s architecture needs to drastically constrain its search space by a bias that is tuned to the application domain (Mitchell, 1997). No such bias is built into the BM’s architecture in its generality  [^1] The concrete application described in Hinton et al. (1984) (the ‘Encoder Problem’) provides such bias, but is trivial in the sense of explicitly defining the full (and very small) set of sensory patterns., while our model’s architecture explicitly incorporates the bias that the sensory patterns of its application domain, vision, are all coherent composites of a finite set of pattern fragments, a bias that is effective yet captures an essentially infinite number of sensory patterns.
The realization that the BM in its generality would imply unfeasible learning times led to the introduction of Restricted Boltzmann Machines (RBM) (Smolenksy, 1987; Hinton and Salakhutdinov, 2006). RBMs consist of two layers, one visible and one latent, with bidirectional connections between them, but no connections within them. A stack of RBMs is used in (Hinton and Salakhutdinov, 2006) to construct an autoencoder in which RBM learning is used to pre-train the connections between layers in a sequence from bottom (input) to top (the sparse coding layer). Then, the whole encoder-decoder stack is trained end-to-end by backpropagation of error. We discuss the differences between our model and networks trained by backpropagation of error (including RBMs) in the next sub-section.
@@ 5.3 Artificial Neural Networks (ANN)
Artificial neural networks (ANNs), as concerned here, are layered structures with feedforward layer-to-layer connections from input to output, and these connections are trained by back-propagation of error (Rosenblatt, 1962, p 292). Prominent examples are multilayer perceptrons (MLPs) (Prince, 2023), CNNs (Lecun et al., 1998), autoencoders (Hinton and Salakhutdinov, 2006), and transformers (Vaswani et al., 2017). In these systems, individual neurons in one layer represent feature compositions that occur in the layer below. The complexity and variability of features represented by individual neurons grow from layer to layer (as shown in reconstructions, see Zeiler and Fergus (2013)). The hierarchy of layers thus constitutes a hierarchy of nested pattern fragments.
The architectures of both ANNs and our model implement the bias that a virtually infinite set of potential input patterns can be represented in the form of flexible compositions of a finite set of pattern fragments. They differ fundamentally, however, in how these pattern fragments are implemented, how they are combined, and how they are learned.
Beginning with the latter, ANNs select the features to be represented on the basis of their effect on lowering the system’s prediction error, whereas in our model the selection of features to be represented by net fragments is entirely a matter of the statistics of input patterns (as is, incidentally, the case with the RBM pretraining of layer connections in Hinton and Salakhutdinov (2006)).
Regarding implementation, ANNs represent elementary pattern fragments in the first latent layer by feature neurons. In our model, which has only one layer, elementary pattern fragments are represented by minimal net fragments, each composed of a neuron and a number of excitatory connections from other neurons within the same layer. That a neuron remains active under attenuation indicates that it has recognized the pattern formed by these other also activated neurons. ANNs represent a nested hierarchy of pattern fragments, as higher-level neurons respond to combinations of active lower-level neurons. Our model equally represents a nested hierarchy of pattern fragments within one layer, in this case composed of a neuron, its immediate supporters, the supporters of the supporters, and so on.
The rules for composing larger structures out of smaller pattern fragments correspondingly differ between ANNs and our model. In CNNs, this flexibility is mostly implemented by max pooling, or in transformers by the attention mechanism that selects active neurons in one layer to contribute to the activity of neurons in the next. This freedom of choosing pattern fragments in lower levels is exerted independently by the neurons in a given level, without regard to the mutual compatibility of those selected lower-level patterns. This lack of coherence is even more drastic between ANNs’ high-level feature neurons, which don’t express in any way mutual overlap in terms of constituent sub-patterns (the binding problem, see Greff et al. (2020)), the likely cause of vulnerability of ANNs to adversarial attack (Leadholm and Stringer, 2022). In our model, in contrast, the net fragments that respond to different pattern fragments have to fit together into one consistent net, in which each active neuron is sufficiently supported by other active neurons. This expresses the bias that significant patterns in the visual environment are composed of pattern fragments in a coherent way. We consider this as the solution to the binding problem (Greff et al., 2020) and as an explanation of the Gestalt phenomenon (Wagemans et al., 2012).
@@ 6 Discussion and Outlook
The Cooperative Network Architecture (CNA) introduced here proposes a novel approach to neural representation by constructing structured, composable nets from input patterns. It is based on the vision to overcome the tension that has eternally been plaguing the brain/mind issue and AI, the tension between low-level and high-level, between sub-symbolic and symbolic descriptions (cf. Fodor and Pylyshyn (1988)). Both perspectives have their strengths and their weaknesses (as discussed by von der Malsburg (2023)). While our model preserves global, high-level pattern structure through the composition of net fragments (the symbolic aspect), it also retains sensitivity to local detail via the fine-grained connectivity of its elements (the sub-symbolic aspect). This synthesis of descriptive levels is the basis for perceptual segmentation (as in the Gestalt phenomenon), noise filtering, pattern completion, and for pattern-global decisions such as classification or recognition (see below).
Empirically, we demonstrated that CNA generalizes beyond its training distribution, successfully reconstructing unseen objects and filtering substantial noise. These findings validate CNA’s ability to form holistic, yet flexible, representations. Furthermore, the reuse and recombination of learned sub-patterns provide a strong inductive bias, allowing for compositional generalization from limited data.
Conceptually, our model deviates in essential ways from classical neural models. In contrast to the perceptron (which is currently dominating AI), the structure of patterns is expressed by recurrent and not feedforward connections, the entirety of patterns is expressed as structured nets instead of by structureless individual neurons, and learning of connectivity is based on Hebbian plasticity and input statistics instead of backpropagation of error and human-provided sample (and ground truth) material. In contrast to the associative memory model, patterns have a flexible articulated structure, being composed of generic pattern fragments, instead of monolithic rigidity. From the Boltzmann Machine, our model differs by being tuned to the natural environment by a powerful bias, which permits to capture a virtual infinity of patterns after learning from a finite set of examples.
What is the functional role of nets within intelligent systems? According to a long-standing proposal, invariant object recognition can be realized by structure-sensitive, homeomorphic mappings between networks (von der Malsburg, 1981; Anderson and van Essen, 1987; Olshausen et al., 1995; Wiskott et al., 1997; Arathorn, 2002; Wolfrum et al., 2008; Memisevic and Hinton, 2010; von der Malsburg et al., 2022). This calls for the existence of a pair or a series of layers, the lowest of which is the sensory layer we have modeled here, while higher layers harbor nets of growing invariance or generality. Nets in higher layers can be invariant against transformations such as translation, rotation, or scaling in the image plane, or can be more abstract by leaving out submodalities such as color or texture, or by accommodating distortion. Dynamic mappings that can keep up with rapid changes within the input layer are realized in the cited literature by the rapid selection of connections that connect neurons of equal feature type and that preserve the neighborhood relationships within the connected nets. Homeomorphic mappings can themselves be considered as nets of a second kind, and they also are coherently composed of overlapping map fragments (Anderson and van Essen, 1987; Zhu and von der Malsburg, 2004). Several nets tied together by mappings can be considered as nets on a higher level. The purpose of the model we have presented here is to fill a sore gap in this line of work in providing a clear picture of how nets that are eligible for mapping are dynamically formed.
Urgent future work will have to expand the present model to cope with a full range of natural visual input and a rich range of visual submodalities (texture, color, stereo depth, motion), will have to expand to other modalities (e.g., audition, motor control), and will have to combine the functional elements modeled here and in the cited work into fully functional systems.